0.001 mol of the strong electrolyte M(OH)2 has been dissolved to make a 20 mL  of its saturated solution. Its pH will be(Kw =1×10^-14)
 

Grade:12th pass

1 Answers

Taniya batham
5 years ago
M(OH)2 → M+ 2OH.                                                     .                            0.001/0.02 *2 =0.1 ( concentration  of. OH-  ions).                                                             [OH] = 0.1
pOH = -log [OH]
.         = 1
pH+ pOH = 14
pH=14-1= 13